OpenStudy (anonymous):

2 particles of mass 30g and 40g respectively both travel at a speed of 35m/s in directions at right angles. The 2 particles collide and stick together. Calculate their speed after impact.

OpenStudy (anonymous):

This is moentum promblem \[m_1v_{1i}+m_2v_{2i}=(m_1+m_2)v_f\\ 30(35\hat{i})+40(35\hat{j})=(30+40)v_f\\ v_f=35\hat{i}+35\hat{j} \]

OpenStudy (anonymous):

I assumes m1 travelling along +ve X-axis and m2 along +ve y-axis resultant travels at a speed of \[35\sqrt{2}\]m/s 45° North of +ve X-axis

OpenStudy (anonymous):

I solved it the same way but the answer sheet say it's 25m/s. O_o

OpenStudy (anonymous):

what they got is \[35/\sqrt{2}\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

but wonder how!!!